OpenStudy (anonymous):

a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)?

OpenStudy (anonymous):

(−2, −2) (−3, −1) (−9, −7) (−5, −3)

pooja195 (pooja195):

Use the substituion method plug in equation 2 into equation one \[\huge~\rm~ b-2 − 3b = 4\]

pooja195 (pooja195):

Now solve for b

OpenStudy (anonymous):

wait how do I solve for b

pooja195 (pooja195):

Start by combining like terms

pooja195 (pooja195):

b−2−3b=4 -3+1=?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

Good now we have -2b now we add 2 to both sides 4+2=?

OpenStudy (anonymous):

6

pooja195 (pooja195):

So now we have -2b=6 Whats 6/-2 ?

OpenStudy (anonymous):

8

pooja195 (pooja195):

?

OpenStudy (anonymous):

no wait u divide those right?

pooja195 (pooja195):

Yes

pooja195 (pooja195):

/=divide

OpenStudy (anonymous):

so than it would be -3

pooja195 (pooja195):

Yes now we plug -3 in place of b in equation 2 a=-3-2 -3-2=?

OpenStudy (anonymous):

-5

pooja195 (pooja195):

good so we have b=-3 a=-5 whats the answer?

OpenStudy (anonymous):

it would be D

pooja195 (pooja195):

Good job! :D

OpenStudy (anonymous):

thx for the help i'll fan and metal now ^_^

pooja195 (pooja195):

thanks ^.^